Find an equation of the circle whose diameter has endpoints (-4,6) and (6,-4)
1
Expert's answer
2013-01-08T11:45:03-0500
equation of the circle:
(x-x0)^2+(y-yo)^2=r^2
where x0, y0 - center, r- radius
r=d/2
d^2=(-4-6)^2+(6+4)^2=200
r^2=d^2/4=50
x0=(-4+6)/2=1
y0=(6-4)/2=1
Answer:
(x-1)^2+(y-1)^2=50